Circular No. 5627

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SCHAUMASSE (1992x) S. Nakano, Sumoto, Japan, reports that this comet has been recovered

by T. Seki, Geisei, using a 0.60-m f/3.5 reflector. The comet was nearly stellar with slight diffuseness on Oct. 1. The prediction on MPC 16379 (ephemeris on MPC 20353) requires correction by ΔT = –0.14 day. The prediction with nongravitational terms by Muraoka on Nakano Note No. 557 (ephemeris on page 60 of Comet Handbook 1992, Oriental Astronomical Assn.) requires correction by ΔT = –0.02 day.

1992 UT  α_{2000}  δ_{2000}  m_{1}
Sept.25.72674  4^{h}25^{m}43.^{s}71  +11^{o}08′15′′.8  20
Oct. 1.77188  4 30 07.34  +11 12 50.8  19.2

SUPERNOVA 1992ay IN ANONYMOUS GALAXY M. M. Phillips, Cerro Tololo Interamerican Observatory, reports that a

spectrogram of SN 1992ay (cf. IAUC 5619) obtained with the CTIO 4.0-m telescope on Oct. 1.34 UT shows that this object is a peculiar type-II event. Strong and relatively narrow (∼ 1200 km/s FWHM) H I emission lines are observed, along with weaker emission apparently due to He I (587.6 and 447.1 nm). The H I emission line profiles appear to be composed of a broad base and a separate, narrower core. These emission features lie atop a strong blue continuum. The cores of the Hα and Hβ lines yield a redshift of 0.062.

PERIODIC COMET SWIFT-TUTTLE (1992t) J. DeYoung and R. Schmidt, U.S. Naval Observatory, report that

P/Swift-Tuttle was observed on Sept. 30.0 UT with the 0.6-m reflector (+ CCD + Hα filter) in Washington. A co-addition of seven 90-s images shows a prominent fan-shaped inner coma with jets or extensions of length 22′′ at p.a. 220^{o} and length 19′′ at p.a. 262^{o}.

Total visual magnitude estimates (cf. IAUC 5621): Sept. 29.01 UT, 9.1 (A. Renou, Blaison, France, 0.12-m reflector); 29.50, 8.9 (C. S. Morris, near Mount Wilson, CA, 20×80 binoculars); 30.77, 9.0 (S. Baroni, Milan, Italy, 20×80 binoculars); Oct. 1.22, 8.9 (M. Tombelli, Montelupo, Italy, 0.44-m reflector).

1992 October 2 Daniel W. E. Green